OpenStudy (anonymous):

what is the slope of a línea parallel to the line 5x+3y=7

OpenStudy (nurali):

a parallel line has the same slope. so first put it in y = mx + b form to find the slope (m). subtract the 5x, and then divide by 3. the new equation is y = -5/3x + 7/3 the slope of a parallel line is still -5/3

OpenStudy (math&ing001):

Parallel lines have the same slope, so you only need to find the slope for 5x+3y=7